OpenStudy (anonymous):

8e^x + 4 lnx^3. Find the derivative.

OpenStudy (kc_kennylau):

Let's do it part by part :)

OpenStudy (anonymous):

ok! thanks

OpenStudy (kc_kennylau):

What's the derivative of 8e^x? :)

OpenStudy (anonymous):

8 or just 8e^x ;-/

OpenStudy (kc_kennylau):

It's 8e^x because e^x doesn't change :)

OpenStudy (kc_kennylau):

Note that the derivative of 8x would be 8 though

OpenStudy (anonymous):

would any of the rules apply? like chain rule? no product or quotient because its not multiplication or a fraction, right?

OpenStudy (anonymous):

ok, noted.

OpenStudy (kc_kennylau):

Not any of the rules that you've learnt would apply to e^x, so you only have to memorize it by heart :)

OpenStudy (anonymous):

I meant apply to the original problem?

OpenStudy (kc_kennylau):

Well, you can treat it as (8e^x)' = 8(e^x)' = 8e^x Then you'd have applied the "multiplication by constant" rule

OpenStudy (kc_kennylau):

I mean \(\Large\frac d{dx}8e^x=8\frac d{dx}e^x=8e^x\)

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

what about the other part of the original question?

OpenStudy (kc_kennylau):

Should we proceed to the second part?

OpenStudy (kc_kennylau):

Oh

OpenStudy (kc_kennylau):

Let's proceed to the second part :)

OpenStudy (anonymous):

ok :-)

OpenStudy (kc_kennylau):

What's the derivative of 4lnx^3?

OpenStudy (anonymous):

I'm not positive how to go about finding the derivative for this part

OpenStudy (anonymous):

use power rule?

OpenStudy (kc_kennylau):

Power rule is not applicable to this problem :)

OpenStudy (kc_kennylau):

Don't forget that ln x^n = nlnx

OpenStudy (anonymous):

so 3lnx

OpenStudy (kc_kennylau):

Don't forget the 4 in front of it :)

OpenStudy (anonymous):

multiply? 12lnx... Im probably making stuff up...i tend to create my own math :-(

OpenStudy (kc_kennylau):

Yep that's correct, let's now proceed to the derivative :)

OpenStudy (anonymous):

yay!

OpenStudy (kc_kennylau):

What's the derivative of 12lnx? :)

OpenStudy (anonymous):

would it be 12? because derivative of lnx =1?

OpenStudy (kc_kennylau):

Derivative of lnx isn't 1 :/

OpenStudy (anonymous):

;-(

OpenStudy (anonymous):

I'm not sure then

OpenStudy (kc_kennylau):

It's 1/x, memorize this :)

OpenStudy (anonymous):

so then it would be 12/x?

OpenStudy (kc_kennylau):

Exactly :)

OpenStudy (kc_kennylau):

Now let's add the result of the first part and the result of the second part together :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8e^x + 12/x, right or no?

OpenStudy (kc_kennylau):

Perfect :D

OpenStudy (anonymous):

you are awesome!

OpenStudy (anonymous):

Can i trouble for help with two more? if not, no worries. I just have to run home and then hop back on

OpenStudy (kc_kennylau):

Sure! :)

OpenStudy (anonymous):

Find the derivative ln(ln7x).

OpenStudy (kc_kennylau):

Chain rule is the way to go :)

OpenStudy (anonymous):

I was thinking chain rule but ln is not a function by itself

OpenStudy (kc_kennylau):

ln is a function by itself

OpenStudy (anonymous):

omg..i think my teacher may be wrong then.

OpenStudy (anonymous):

so u=ln7x u'= ln(u)

OpenStudy (kc_kennylau):

Does the apostrophe mean derivative of?

OpenStudy (anonymous):

yes

OpenStudy (kc_kennylau):

Then how would u' be equal to ln(u)

OpenStudy (kc_kennylau):

y=ln(u) though

OpenStudy (anonymous):

the problem says find f'x for f(x)=ln(ln7x)

OpenStudy (kc_kennylau):

Okay I mean f(x)=ln(u)

OpenStudy (kc_kennylau):

Where u=ln(7x)

OpenStudy (kc_kennylau):

Now find f'(x) :)

OpenStudy (anonymous):

Ok, I'm thinking f'(x) of ln7x is 7/x, right or wrong?

OpenStudy (kc_kennylau):

Wrong

OpenStudy (anonymous):

wait, it would be 1/x

OpenStudy (kc_kennylau):

1. There's no such thing as f'(x) of ln7x because f(x) is already defined to be ln(ln(7x))

OpenStudy (kc_kennylau):

2. Yep it's 1/x

OpenStudy (anonymous):

are you a math teacher? :-)

OpenStudy (kc_kennylau):

No, I'm Form 3 (grade 9 in US educational system)

OpenStudy (kc_kennylau):

I'm probably younger than you lol

OpenStudy (anonymous):

no way! so youre like how old?

OpenStudy (kc_kennylau):

14 years old :)

OpenStudy (anonymous):

lol! i feel like such an idiot now. yowza! i am in pre-calc for college. *puts head down in disgrace*

OpenStudy (kc_kennylau):

xD

OpenStudy (anonymous):

ok, back to original problem...

OpenStudy (kc_kennylau):

Find f'(x) :)

OpenStudy (kc_kennylau):

\[\frac d{dx}\ln(\ln(7x))=\frac d{d\ln(7x)}\ln(\ln(7x))\times\frac d{dx}\ln(7x)\]

OpenStudy (anonymous):

that looks like a foreign language to me... sigh ok, let me try

OpenStudy (anonymous):

the ds cancel out, and youre left with 1/ln(7x)

OpenStudy (kc_kennylau):

No the ds don't cancel out

OpenStudy (anonymous):

darnit

OpenStudy (kc_kennylau):

\(\dfrac d{dx}g(x)\) means \(g'(x)\)

OpenStudy (kc_kennylau):

The ds are just notation

OpenStudy (anonymous):

oh i see i do remember seeing that

OpenStudy (anonymous):

1/ln(7x) and im stuck now

OpenStudy (kc_kennylau):

\[[f(g(x)]'=f'(g(x))\times g'(x)\]

OpenStudy (kc_kennylau):

u(x)=ln(7x) v(x)=ln(x) f(x)=v(u(x)

OpenStudy (kc_kennylau):

f'(x) = [v(u(x))]' = v'(u(x)) * u'(x)

OpenStudy (anonymous):

so 1/x (ln(7x)) * u'(x).....I know im prolly not getting it :-( u'(x)=1/x

OpenStudy (kc_kennylau):

For example, \(\sin(5x)'=\cos(5x)\times5\)

OpenStudy (kc_kennylau):

For example, \(\cos(\sqrt x)'=-\sin(\sqrt x)\times\dfrac1{2\sqrt x}\)

OpenStudy (anonymous):

1/ln(7x) * (1/7x)= 1/7xlnx, correct or no?

OpenStudy (kc_kennylau):

It's not 1/7x

OpenStudy (anonymous):

right you said that, excuse my brain! its 1/ln(7x) * (1/x)= 1/7xlnx

OpenStudy (kc_kennylau):

Again, it's not 1/7x

OpenStudy (anonymous):

:-/ ok, give me sec

OpenStudy (kc_kennylau):

The part to the left of the equal sign is correct

OpenStudy (anonymous):

ok, so does it get simplified further or no? 1/ln(7x) * (1/7x), can I leave it as is?

OpenStudy (anonymous):

i meant 1/x* sorry

OpenStudy (kc_kennylau):

\[\frac1{\ln(7x)}\times\frac1x=?\]

OpenStudy (anonymous):

1/ ln (7x) (x), im not sure if i should square x

OpenStudy (kc_kennylau):

Perfect

OpenStudy (anonymous):

yay! so f'(x) of f(x) =ln(ln7x)= 1/ ln (7x) (x)